OpenStudy (anonymous):

Simplify -8/3 / -2/6

OpenStudy (debbieg):

Well, first, division by a fraction is equivalent to multiplication by the reciprocal. So you have: \[\large \frac{ \frac{ -8 }{ 3 }}{\frac{ -2 }{ 6 } }=\frac{ -8 }{ 3 }\times \frac{ 6 }{ -2 }\]

OpenStudy (debbieg):

Then, you should always reduce common factors first, before multiplying. Once you do that, (the negative signs cancel out; the 8 and 2 reduce, and 6 and 3 reduce), what do you get for an answer?

OpenStudy (anonymous):

8?

OpenStudy (debbieg):

(Or even, just reduce the 6 and the 2, cancel the negatives, and then one more cancellation will quickly give you the resulting answer.)

OpenStudy (debbieg):

Right!

OpenStudy (anonymous):

Thank you I appreciate it!

OpenStudy (debbieg):

You're very welcome. :)